Please answer all parts, Thanks!P3-24 (similar to) Waiting period. Jamal is waiting to be a millionaire. He wants to know how long he must wait if a. he invests $27,896.91 at 21% today? b, he invests $48,188.43 at 13% today? c, he invests $157,694.02 at 10% today? d. he invests $274,559.09 at 5% today? a. How long will Jamal have to wait to become a millionaire if he invests $27,896.91 at 21% today? years (Round to the nearest whole number.)
Answer #1

a.

FV = $1,000,000

PV = $27,896.91

Interest rate = 21%

Using TVM calculation,

FV = PV(1+r)t

T = 18.74 years

b.

FV = $1,000,000

PV = $48,188.43

Interest rate = 12%

Using TVM calculation,

FV = PV(1+r)t

T = 24.65 years

c.

FV = $1,000,000

PV = $157,694.02

Interest rate = 10%

Using TVM calculation,

FV = PV(1+r)t

T = 19.4 years

d.

FV = $1,000,000

PV = $274,559.09

Interest rate = 5%

Using TVM calculation,

FV = PV(1+r)t

T = 26.64 years